• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 16:00:55 on March 4,2007

    Entry number 194721

    keyword=Shift summary

    8:00 doing delta scan with sieve slits out, -2%;
    9:17 finished delta scan, called MCC to start pass change, ramping down HRSR and cycling HRSL Q2/Q3;
    9:36 left Q3 quenched after typing in the new momentum setting;
    12:30 Hall back to beam permit, during controlled access the HRSR polarity was changed to POSITIVE and the angle was moved to about 25 degrees; The techs also fixed the HRSL Q3 quench (had to increase the Q3 helium flow to prevent multiple quench. Had problem with left dipole too but was fixed at the same access).

    12:45 Pulsed beam are expected in the hall for tuning;
    12:58 HRSR is now at 54.10 degrees with floor mark 54.0 and vernier -17mm;
    12:59 HRSL fields are now fully recovered; Set HRSR momentum to 0.989 GeV/c;

    16:00 MCC said they will pulsed (at 2pm) then CW beam (3pm) for tuning, but we haven't seen much beam in the hall yet. Right dipole still ramping up at the end of shift.